OpenStudy (anonymous):

Find the value of Z in the similar figures (round to the nearest tenth) A. 12 B. 6 C. 13.3 D. 7.3 Give me a second, I gota draw it:)

OpenStudy (anonymous):



OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

8/z = 4/6 .... .... z = ??

OpenStudy (anonymous):

z=12

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

good job